OpenStudy (anonymous):

Choose the equation of the line passing through the point (2, 6) and parallel to y = -3x - 4. A. y = 3x - 12 B. y = 3x - 12 C. y = -3x + 12 D. y = 3x

OpenStudy (anonymous):

guess man guess!!

OpenStudy (triciaal):

parallel lines have the same slope slope is -3 using the point (2,6) find the intercept 6 = -3(2) + b 6 = -6 + b ; b = 12 line y = -3x + 12 answer choice C